OpenStudy (jazzyfa30):

Match the equation with its solution. x4 -13x2 + 36 =0 x4 - 11x2 + 24 =0 x - 7 \sqrt{x} =-12 x - 5 \sqrt{x} +6=0 x4 + x2 =12

OpenStudy (whpalmer4):

don't see any solutions...

OpenStudy (whpalmer4):

Well, here's a hint: anything with an \(x^4\) in it here will have 4 solutions. Everything else will have 2 solutions.

OpenStudy (whpalmer4):

You can try factoring, or try out the possible answers in an equation and see if they work.

OpenStudy (whpalmer4):

Do you know the rational root/zero theorem?

OpenStudy (jazzyfa30):

no

OpenStudy (whpalmer4):

Okay, well, I won't suggest you use it then :-) Let's work through the first one as an example: \[x^4-13x^2+36=0\]Could you factor that if it was\[x^2-13x+36=0\]?

OpenStudy (whpalmer4):

you need to find a pair of numbers that multiply to 36, and add to -13

OpenStudy (jazzyfa30):

-4 and 9

OpenStudy (whpalmer4):

how about -4 and -9? That would be \[ (x-4)(x-9) = x^2-9x-4x+36 = x^2-13+36\] Well, the only difference between that and what you have is the exponents are twice as big here. It factors the same way: \[(x^2-4)(x^2-9) = x^4-9x^2-4x^2+36 = x^4-13x^2+36\] But wait, notice that both of the factors can be factored again, as the difference of squares: \[x^2-4 = (x-2)(x+2)\]\[x^2-9 = (x-3)(x+3)\]That means that \[(x-2)(x+2)(x-3)(x+3) = x^4-13x^2+36=0\]And what are the solutions to all of those product terms?

OpenStudy (whpalmer4):

(-4 and 9 don't work, has to be -4 and -9)

OpenStudy (jazzyfa30):

ok

OpenStudy (whpalmer4):

x-2 = 0, x = x+2 = 0, x = x-3 = 0, x = x+3 = 0, x =

OpenStudy (jazzyfa30):

x=2 x=-2 x=3 x=-3

OpenStudy (whpalmer4):

Right! there's a shorthand sometimes used \(x = \pm 2, \pm 3\) which means the same thing as your list

OpenStudy (whpalmer4):

and they use a word version of it in the answer list

OpenStudy (jazzyfa30):

umm the answer is c

OpenStudy (whpalmer4):

why would you say that?

OpenStudy (jazzyfa30):

ummmmmm idk its d

OpenStudy (whpalmer4):

Come on, you're guessing, and there's no need to guess when you have the answer right in front of you. Only one of those answer choices gives you -2 +2 -3 and +3

OpenStudy (whpalmer4):

do you see any square root signs in our answer?

OpenStudy (jazzyfa30):

ohhhhh the answer is E and no

OpenStudy (whpalmer4):

that's better :-)

OpenStudy (whpalmer4):

okay, I'll help you with one more of the other kind: \[x-7\sqrt{x} = -12\]Subtract x from both sides \[-7\sqrt{x} = -12-x\]Square both sides \[(-7\sqrt{x})^2 = (-12-x)(-12-x)\]\[49x^2 = 144+12x+12x+x^2\]\[49x^2=144+24x+x^2\]Subtract \(49x^2\) from both sides \[0=144+24x-48x^2\]Factor or complete square or use quadratic formula

OpenStudy (jazzyfa30):

???????

OpenStudy (whpalmer4):

Whoops, sorry, that should be \[49x = 144+24x + x^2\] \[0=144-25x+x^2\]

OpenStudy (whpalmer4):

What's your question?

OpenStudy (whpalmer4):

\[x^2-25x+144=0\] Two numbers that multiply to 144 and sum to -25: -9, -16 \[(x-9)(x-16)=0\]\[x=\]

OpenStudy (whpalmer4):

I have to go, good luck with the others. You should be able to do them by following these examples.